OpenStudy (anonymous):

Which of the following would eliminate the variable on the left side of the given equation? 36 - 19w = -6w + 41 add -19w add 6w add 19w add -36

OpenStudy (anonymous):

@dtuniyants

OpenStudy (anonymous):

gsh add 19w maybe?

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

Simplifying 36 + -19w = -6x + 41 Reorder the terms: 36 + -19w = 41 + -6x Solving 36 + -19w = 41 + -6x Solving for variable 'w'. Move all terms containing w to the left, all other terms to the right. Add '-36' to each side of the equation. 36 + -36 + -19w = 41 + -36 + -6x Combine like terms: 36 + -36 = 0 0 + -19w = 41 + -36 + -6x -19w = 41 + -36 + -6x Combine like terms: 41 + -36 = 5 -19w = 5 + -6x Divide each side by '-19'. w = -0.2631578947 + 0.3157894737x Simplifying w = -0.2631578947 + 0.3157894737x